OpenStudy (anonymous):

4(9x - 12) = 3(16x - 12)

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

4(9x - 12) = 3(16x - 12) 36x - 48 = 48x - 36 -12 = 12x x = -1

OpenStudy (anonymous):

this is interesting graph y=36x-48, and y=48x-36, the intersection is the solution will y=ax-b and y=bx-a always have x=-1?

OpenStudy (anonymous):

36x - 48 = 48x- 36 +48 +48 36x = 48x+12 -48x -48x -12x = 12 divide by -12 -1x = 1 x = -1